Bugs in Computers & Typesetting, 1991 1

This is a list of all corrections made to Computers & Typesetting, Volumes A, B, C, and D, between 1 January 1991 and 15 March 1992. Corrections made to the softcover version of The TEXbook are the same as corrections to Volume A. Corrections to the softcover version of The book are the same as corrections to Volume C. Some of the corrections below have already been made in reprintings of the books. Changes to Volume B refer to the fourth printing (1991), which differs markedly from earlier printings because it includes all the revisions for TEX3.0. Changes to Volume D refer to the third printing (1991), which differs markedly from earlier printings because it includes all the revisions for 2.0. Changes to the mini-indexes and master indexes of Volumes B and D are not shown here unless they are not obviously derivable from what has been shown. Dozens of changes, too many to list here, have been made to Volume E because of recent upgrades to the Computer Modern font source files. Those changes, which affect only the digitization at low resolution and the appearance of lowercase delta and some characters in the math symbols fonts (but not the TFM files), are documented at the end of file cm85.bug.

Page A96, lines 9–11 (9/18/91)

Some German words traditionally change their spelling when they are split between lines. For example, ‘backen’ becomes ‘bak-ken’ and ‘Bettuch’ becomes ‘Betttuch’. How can you instruct TEX to produce such effects?

Page A178, line 17 (11/19/91)

If you say ‘\phantom{⟨subformula⟩}’ in any formula, plain TEX will do its

Page A286, bottom two lines and continuing into A287 (11/21/91)

stands for zero or more ⟨assignment⟩ commands other than \setbox. If the assignments are not followed by a ⟨character⟩, where ⟨character⟩ stands for any of the commands just discussed in the previous paragraph, TEX treats \accent as if it were \char, except that the space factor is set to 1000. Otherwise the character that follows the assignment is accented by the character that corresponds to the ⟨8-bit number⟩. (The purpose of the intervening assignments is to allow the accenter and accentee to be in different fonts.) If the accent must be moved up or down, it is put into an hbox that is raised or lowered. Then the accent is effectively superposed on the character by means of kerns, in such a way that the width of the accent does not influence the width of the resulting horizontal list. Finally, TEX sets \spacefactor=1000.

Page A291, lines 6–8 (11/21/91)

‘}’ may be followed by optional ⟨assignment⟩ commands other than \setbox, after which ‘$$’ must conclude the display. TEX will insert the \abovedisplayskip and \belowdisplayskip glue before and after the result of the alignment.

Page A293, line 14 (9/18/91)

explained in Appendix G. TEX scans ⟨one optional space⟩ after completing a displayed formula; this is usually the implicit space at the end of a line in the input file.

Page A311, bottom four lines (9/18/91)

12.7. 1000, except: 999 after O, B, S, D, and J; 1250 after the comma; 3000 after the exclamation point, the right-quote marks, and the periods. If a period had come just after the B (i.e., if the text had said ‘B. Sally’), the space factor after that period would have been 1000, not 3000.

Page A314, lines 16–18 from the bottom (1/10/92)

14.8. ba\ck/en and Be\ttt/uch, where the macros \ck/ and \ttt/ are defined by

        \def\ck/{\discretionary{k-}{k}{ck}}
        \def\ttt/{tt\discretionary{-}{t}{}}

Page A354, line 8 (9/18/91)

\def\multispan#1{\omit\mscount=#1\relax\loop\ifnum\mscount>1 \sp@n\repeat}

Page A356, line 11 from the bottom (9/23/91)

  \else{\ooalign{\unhbox0\crcr\hidewidth\char'30\hidewidth}}\fi}

Page A358, line 8 from the bottom (9/18/91)

\mathchardef\mapstochar="3237 \def\mapsto{\mapstochar\rightarrow}

Page A359, line 13 (11/4/91)

\def\overrightarrow#1{\vbox{\m@th\ialign{##\crcr

Page A359, line 16 (11/4/91)

\def\overleftarrow#1{\vbox{\m@th\ialign{##\crcr

Page A359, line 19 (11/4/91)

\def\overbrace#1{\mathop{\vbox{\m@th\ialign{##\crcr\noalign{\kern3pt}

Page A359, line 22 (11/4/91)

\def\underbrace#1{\mathop{\vtop{\m@th\ialign{##\crcr

Page A359, lines 7–14 from the bottom (1/11/92)

\def\lgroup{\delimiter"462833A }      \def\rgroup{\delimiter"562933B }
\def\lmoustache{\delimiter"437A340 }  \def\rmoustache{\delimiter"537B341 }
\def\uparrow{\delimiter"3222378 }     \def\Uparrow{\delimiter"322A37E }
\def\downarrow{\delimiter"3223379 }   \def\Downarrow{\delimiter"322B37F }
\def\updownarrow{\delimiter"326C33F } \def\arrowvert{\delimiter"026A33C }
\def\Updownarrow{\delimiter"326D377 } \def\Arrowvert{\delimiter"026B33D }
\def\vert{\delimiter"026A30C }        \def\Vert{\delimiter"026B30D }
\def\backslash{\delimiter"026E30F }   \def\bracevert{\delimiter"077C33E }

Page A360, line 13 (11/19/91)

\phantom, \smash, \root, and other operations. (Actually \phantom and \smash are not perfect: They assume that the current style is uncramped.)

Page A360, line 2 from the bottom (11/4/91)

\def\c@ncel#1#2{\m@th\ooalign{$\hfil#1\mkern1mu/\hfil$\crcr$#1#2$}}

Page A361, top line (11/4/91)

\def\rlh@#1{\vcenter{\m@th\hbox{\ooalign{\raise2pt

Page A364, line 5 from the bottom (11/4/91)

\def\fmtname{plain}\def\fmtversion{3.141}

Page A377, the bottom 17 lines (9/18/91)

story: Macro \stest decides whether or not a given token list register begins with a ⟨space token⟩ as defined in Chapter 24. If so, the macro decides whether the token is explicit and/or funny and/or active.

\newif\ifspace \newif\iffunny \newif\ifexplicit \newif\ifactive
\def\stest#1{\funnyfalse \expandafter\s\the#1! \stest}
\def\s{\global\explicitfalse \global\activefalse \futurelet\next\ss}
\def\ss{\ifcat\noexpand\next\stoken\let\nxt\sx\else\let\nxt\ns\fi\nxt}
\def\sx{\spacetrue\ifx\next\stoken\let\nxt\sss\else\let\nxt=\ssss\fi\nxt}
\long\def\sss#1 #2\stest{\def\next{#1}%
  \ifx\next\empty \global\explicittrue \else\testactive#1\s\fi}
\long\def\ssss#1#2\stest{\funnytrue{\escapechar=\if*#1`?\else`*\fi\relax
  \if#1\string#1\uccode`#1=`~ % we assume that ~ is an active character
    \uppercase{\ifcat\noexpand#1}\noexpand~\global\activetrue
    \else\global\explicittrue\fi
  \else\testactive#1\s\fi}}
\long\def\ns#1\stest{\spacefalse}
\long\def\testactive#1#2\s{\expandafter\tact\string#1\s\tact}
\long\def\tact#1#2\tact{\def\next{#2}\ifx\next\xs\global\activetrue
  \else\ifx\next\empty \global\activetrue\fi\fi} \def\xs{\s}

Page A444, lines 15–26 (3/26/91)

14. If the current item is an Ord atom, go directly to Rule 17 unless all of the following are true: The nucleus is a symbol; the subscript and superscript are both empty; the very next item in the math list is an atom of type Ord, Op, Bin, Rel, Open, Close, or Punct; and the nucleus of the next item is a symbol whose family is the same as the family in the present Ord atom. In such cases the present symbol is marked as a text symbol. If the font information shows a ligature between this symbol and the following one, using the specified family and the current size, then insert the ligature character and continue as specified by the font; in this process, two characters may collapse into a single Ord text symbol, and/or new Ord text characters may appear. If the font information shows a kern between the current symbol and the next, insert a kern item following the current atom. As soon as an Ord atom has been fully processed for ligatures and kerns, go to Rule 17.

Page A446, lines 5 and 6 from the bottom (1/13/92)

are used to change the current style just as in the first pass, so that both passes have the same value of C when they work on any particular atom.

Page A447, in the parameter usage table (1/13/92)

[Delete the entry for ‘σ_{2}’; the entry for ‘σ_{17}’ moves down to the bottom of the left column.]

Page A447, line 2 after the parameter usage table (1/13/92)

to parameters in arbitrary families: Rule 17 uses \fontdimen parameter 2 (space) to de-

Page A467, entry for \hss (9/18/91)

*\hss, 71–72, 82–83, 233, 283, 285, 290, 442.

Page A467, new subentry under hyphenation (9/18/91)

suppressing, 93, 414, 424, 454.

Page A476, right column (11/21/91)

*\setbox, 66–67, 77, 81, \underline{120}, 276, 279, 286,

291, 386–392.

Page B2, line 10 from the bottom (1/11/92)

define banner ≡ ´ This␣is␣TeX,␣Version␣3.141´ { printed when TEX starts }

Page B18, lines 21 and 22 (10/12/91)

must have an xchr equivalent in the local character set. (This restriction applies only to preloaded strings, not to those generated dynamically by the user.)

Page B26, new line before fourth line from bottom (1/24/92)

nl : integer ; { new-line character to restore }

Page B26, bottom line and top 3 lines of B27 (1/24/92)

else begin if selector > pseudo then begin print char (s); return; { internal strings are not expanded } end;

if (⟨ Character s is the current new-line character 244 ⟩) then

if selector < pseudo then

begin print ln ; return; end; nl ← new line char ; new line char ← −1; { temporarily disable new-line character } j ← str start [s]; while j < str start [s + 1] do

begin print char (so(str pool [j])); incr (j); end; new line char ← nl ; return; end;

Page B27, lines 9 and 10 (9/19/91)

60. Control sequence names, file names, and strings constructed with \string might contain ASCII code values that can’t be printed using print char . Therefore we use slow print for them:

Page B27, lines 13–26 (1/24/92)

var j: pool pointer ; { current character code position } begin if (s ≥ str ptr ) ∨ (s < 256) then print (s) else begin j ← str start [s];

while j < str start [s + 1] do begin print (so(str pool [j])); incr (j); end;

end; end;

Page B28, line 8 (9/19/91)

else begin slow print (format ident ); print ln ;

Page B33, line 3 (1/11/92)

recursively. A similar interlock is provided by set box allowed .

Page B33, new line to come after line 14 (1/11/92)

set box allowed : boolean ; { is it safe to do a \setbox assignment? }

Page B33, new line to come after line 20 (1/11/92)

set box allowed ← true;

Page B36, line 12 (9/19/91)

begin print nl ("You␣want␣to␣edit␣file␣"); slow print (input stack [base ptr ].name field );

Page B46, lines 9 and 10 (5/24/91) arithmetic; see TUGboat 3,1 (March 1982), 10–27. (But the routines cited there must be modified to allow negative glue ratios.)

Page B47, lines 2 and 3 (5/24/91) structures on a memory word , which contains either a (signed) integer, possibly scaled, or a (signed) glue ratio, or a small number of fields that are one half or one quarter of the size used

Page B177, lines 10 and 11 (9/19/91)

begin print err ("Bad␣mathchar");

help2 ("A␣mathchar␣number␣must␣be␣between␣0␣and␣32767.")

Page B196, new lines after line 11 (1/13/92)

if align state < 1000000 then { unmatched ‘}’ aborts the line } begin repeat get token ; until cur tok = 0; align state ← 1000000; goto done; end;

Page B208, line 21 (9/19/91)

begin slow print (a); slow print (n); slow print (e);

Page B214, line 14 (9/19/91)

begin wlog (banner ); slow print (format ident ); print ("␣␣"); print int (day ); print char ("␣");

Page B214, line 2 from the bottom (9/19/91)

print char ("("); incr (open parens ); slow print (name); update terminal ; state ← new line;

Page B234, line 22 (9/19/91)

print ("␣in␣font␣"); slow print (font name[f]); print char ("!"); end diagnostic(false);

Page B267, lines 7 and 8 (9/19/91)

print nl ("Output␣written␣on␣"); slow print (output file name); print ("␣("); print int (total pages ); print ("␣page");

Page B296, new lines after line 8 of section 716 (1/11/92)

if f < 0 then begin decr (n); f ← f + ´ 200000 ; end;

Page B297, new lines after line 7 of section 717 (1/11/92)

if f < 0 then begin decr (n); f ← f + ´ 200000 ; end;

Page B348, bottom two lines (1/3/92)

Up to three passes might be made through the paragraph in an attempt to find at least one set of feasible breakpoints. On the first pass, we have threshold = pretolerance and second pass =

Page B364, line 20 (1/3/92)

863. The ‘loop’ in the following code is performed at most thrice per call of line break , since

Page B377, insert new line after line 12 (9/19/91)

hyf bchar : halfword ; { boundary character after c_{n} }

Page B378, line 12 from the bottom (9/19/91)

hyf bchar ← character (s); c ← qo(hyf bchar );

Page B378, line 9 from the bottom (1/10/92)

hb ← s; incr (hn ); hu [hn ] ← c; hc[hn ] ← lc code(c); hyf bchar ← non char ;

Page B378, line 5 from the bottom (9/19/91)

else if (type(s) = kern node) ∧ (subtype(s) = normal ) then hb ← s else goto done3 ;

Page B379, line 6 (9/19/91)

j ← hn ; q ← lig ptr (s); if q > null then hyf bchar ← character (q);

Page B379, new line between lines 14 and 15 (1/10/92)

if odd (subtype(s)) then hyf bchar ← font bchar [hf ] else hyf bchar ← non char ;

Page B379, line 19 (9/19/91)

if hn < l hyf + r hyf then goto done1 ; { l hyf and r hyf are always ≥ 1 }

Page B380, lines 9–11 from the bottom reduce to a single line (1/10/92)

q ← link (hb); link (hb) ← null ; r ← link (ha ); link (ha ) ← null ; bchar ← hyf bchar ;

Page B436, lines 9 and 10 (3/15/92)

cur r = \{ \begin{matrix}character (lig stack ), & if lig stack > null ; \\ font bchar [cur font ], & otherwise; \end{matrix}

except when character (lig stack ) = font false bchar [cur font ]. Several additional global variables are needed.

Page B438, line 13 from the bottom (3/15/92)

cur q ← tail ; cur l ← character (lig stack );

Page B507, line 6 of section 1241 (1/11/92)

scan optional equals ; if set box allowed then scan box (box flag + n) else begin print err ("Improper␣"); print esc("setbox");

  help2 ("Sorry,␣\setbox␣is␣not␣allowed␣after␣\halign␣in␣a␣display,")
  ("or␣between␣\accent␣and␣an␣accented␣character."); error ;

end;

Page B511, new line inserted after line 3 (1/24/92)

flushable string : str number ; { string not yet referenced }

Page B512, new line inserted after line 3 of section 1260 (1/24/92)

flushable string ← str ptr − 1;

Page B512, the former line 6 of section 1260 (1/24/92)

begin if cur name = flushable string then

begin flush string ; cur name ← font name[f]; end; if s > 0 then

Page B512, line 10 from the bottom (9/19/91)

set font : begin print ("select␣font␣"); slow print (font name[chr code]);

Page B514, line 9 (1/11/92)

set box allowed ← false; prefixed command ; set box allowed ← true;

Page B515, line 19 (9/19/91)

slow print (s); update terminal ;

Page B516, line 2 (9/19/91)

begin print err (""); slow print (s);

Page B531, lines 19 and 20 (9/19/91)

print nl ("Beginning␣to␣dump␣on␣file␣"; slow print (w make name string (fmt file)); flush string ; print nl (""); slow print (format ident )

Page B533, line 29 (9/19/91)

begin print nl ("Transcript␣written␣on␣"); slow print (log name); print char (".");

Page B538, line 13 (9/19/91)

10: slow print (n);

Page B577, left column (12/23/91)

[Add 798 to the index entries for ‘system dependencies’.]

Page C262, line 15 (3/26/91)

string base_name, base_version; base_name="plain"; base_version="2.7";

Page C271, line 17 from the bottom (3/26/91)

  currentpen_path shifted (z.t_) withpen penspeck enddef;

Page C347, Bront”e entry (1/29/91)

[The accent was clobbered; her name should, of course, be Brontë. Fix the entries for Dürer, Möbius, and Stravinsky in the same way.]

Page C348, left column (1/11/92)

compound statement, \underline{155}, 217.

Page C353, right column (1/11/92)

*numeric, 55, \underline{56}, 65, 88.

Page C354, miscellaneous entries in both columns (1/11/92)

*openwindow, \underline{191}–\underline{193}, 220, 277, 312–313.

*or, 65, \underline{170}, 210, 237, 288–289.
*pair, 55, \underline{56}, 65.
*path, 55, \underline{56}, 171.
*pen, 55, \underline{56}, 65, 170.
*picture, 55, \underline{56}, 114.

Page C356, right column (1/11/92)

*string, 55, \underline{56}, 69.

Page C357, right column (1/11/92)

*transform, 55, \underline{56}, 57, 141–143, 160, 266.

Page D2, last line of section 2 (1/24/92)

define banner ≡ ´ This␣is␣METAFONT,␣Version␣2.71´ { printed when starts }

Page D102, line 15 from the bottom (11/1/91)

Then eq type(h(x)) = tag token and equiv (h(x)) = p, where p is a two-word value node with

Page D188, lines 16 and 17 (1/24/92)

errors. Our subroutines also obey the identity t[a, b] + t[b, a] = a + b.

Page D190, new copy before bottom four lines (1/24/92)

if x coord (r) < x coord (pp) then x coord (r) ← x coord (pp) else if x coord (r) > dest x then x coord (r) ← dest x ; if left x (r) > x coord (r) then

begin left x (r) ← x coord (r); if right x (pp) > x coord (r) then right x (pp) ← x coord (r); end; if right x (r) < x coord (r) then

begin right x (r) ← x coord (r); if left x (qq ) < x coord (r) then left x (qq ) ← x coord (r); end;

Page D191, new copy before bottom two lines of section 416 (1/24/92)

if x coord (s) < x coord (r) then x coord (s) ← x coord (r) else if x coord (s) > dest x then x coord (s) ← dest x ; if left x (s) > x coord (s) then

begin left x (s) ← x coord (s); if right x (r) > x coord (s) then right x (r) ← x coord (s); end; if right x (s) < x coord (s) then

begin right x (s) ← x coord (s); if left x (qq ) < x coord (s) then left x (qq ) ← x coord (s); end;

Page D194, lines 4 and 5 (1/24/92)

[Delete those two lines; I no longer believe that the assertion has been proved (although it might be true).]

Page D194, lines 7–13 of section 424 (1/24/92)

if y coord (r) < y coord (p) then y coord (r) ← y coord (p) else if y coord (r) > dest y then y coord (r) ← dest y ; if x coord (p) + y coord (r) > dest x + dest y then y coord (r) ← dest x + dest y − x coord (p); if left y (r) > y coord (r) then

begin left y (r) ← y coord (r); if right y (p) > y coord (r) then right y (p) ← y coord (r); end; if right y (r) < y coord (r) then

begin right y (r) ← y coord (r); if left y (q) < y coord (r) then left y (q) ← y coord (r); end;

Page D194, lines 8–11 from the bottom (1/24/92)

if right y (r) < y coord (r) then begin right y (r) ← y coord (r); if left y (q) < y coord (r) then left y (q) ← y coord (r); end;

Page D195, lines 3–9 of section 425 (1/24/92)

if y coord (s) < y coord (r) then y coord (s) ← y coord (r) else if y coord (s) > dest y then y coord (s) ← dest y ; if x coord (r) + y coord (s) > dest x + dest y then y coord (s) ← dest x + dest y − x coord (r); if left y (s) > y coord (s) then

begin left y (s) ← y coord (s); if right y (r) > y coord (s) then right y (r) ← y coord (s); end; if right y (s) < y coord (s) then

begin right y (s) ← y coord (s); if left y (q) < y coord (s) then left y (q) ← y coord (s); end;

Page D195, lines 3–7 from the bottom if section 425 (1/24/92)

if right y (s) < y coord (s) then begin right y (s) ← y coord (s); if left y (q) < y coord (s) then left y (q) ← y coord (s); end;

Page D289, lines 9 and 10 (11/1/91)

p ← dep list (p); r ← inf val ;

repeat if value(info(p)) ≥ value(r) then

Page D486, line 18 (11/1/91)

The label loc and label char arrays have been set up to record all the starting addresses; we have